OpenStudy (lolipop_drum312):

What is the solution to the linear inequality −1/4x<8 ? _ Enter your answer in the box.

OpenStudy (lolipop_drum312):

@civicsiscool44

OpenStudy (lolipop_drum312):

um ok so do we simplify them

OpenStudy (lolipop_drum312):

?

OpenStudy (civicsiscool44):

hold on a sec

OpenStudy (lolipop_drum312):

but i have to put a number

OpenStudy (civicsiscool44):

\[\frac{- 1 }{ 4 } x <8\]

OpenStudy (civicsiscool44):

Is it written like that

OpenStudy (lolipop_drum312):

no its written like how its written on top

OpenStudy (lolipop_drum312):

or do you mean its x<8?

OpenStudy (lolipop_drum312):

i mean yes

OpenStudy (civicsiscool44):

Thats what I meant for how it is written

OpenStudy (lolipop_drum312):

@3mar

OpenStudy (3mar):

Well, I am here.

OpenStudy (lolipop_drum312):

i ahve to solve those

OpenStudy (lolipop_drum312):

and idk how

OpenStudy (3mar):

How can I help you?

OpenStudy (lolipop_drum312):

it witht he question up top do i add or or somthing

OpenStudy (lolipop_drum312):

or simplify

OpenStudy (3mar):

from the top?

OpenStudy (lolipop_drum312):

yes the q on top

OpenStudy (3mar):

Ok Let me see

OpenStudy (3mar):

\[\frac{ - 1}{ 4x }<8\] THis is the form?

OpenStudy (lolipop_drum312):

yes

OpenStudy (lolipop_drum312):

wait the the x is on the side of the line

OpenStudy (3mar):

Ok What you think how to start?

OpenStudy (3mar):

\[\frac{ -1 }{ 4 }x<8\]

OpenStudy (lolipop_drum312):

yes..and um idk i think get rid of x idk

OpenStudy (3mar):

Not quite. Let's firstly separate x in one side and the others in the other side. ok?

OpenStudy (3mar):

Can you?

OpenStudy (civicsiscool44):

\[x>-32\]

OpenStudy (lolipop_drum312):

hm? is that the answer or only the beginning

OpenStudy (civicsiscool44):

Thats your answer sorry i did the question wrong

OpenStudy (lolipop_drum312):

it fine

OpenStudy (lolipop_drum312):

i'm gonna do a recheck in another panel to see if my answer are right

OpenStudy (lolipop_drum312):

becuzz i do not wanna take this test again

OpenStudy (3mar):

Do you follow?

OpenStudy (3mar):

Please post a new one and let's start the answer.